OpenStudy (anonymous):

write the ratio as a fraction in lowest terms. compare in inches 4 feet to 31 inches

OpenStudy (anonymous):

first you gotta get them both numbers in the same measure ment

OpenStudy (anonymous):

I think 12 inches is a foot

OpenStudy (anonymous):

4*12=48

OpenStudy (anonymous):

48/31

OpenStudy (anonymous):

oh whoops the opposite

OpenStudy (anonymous):

31 over 48 i would believe

OpenStudy (anonymous):

31 *4=124

OpenStudy (anonymous):

whoops that is also wrong

OpenStudy (anonymous):

no because its says in simplest form!

OpenStudy (anonymous):

or lowest form (same thing)

OpenStudy (anonymous):

31/48 is wrong=/

OpenStudy (anonymous):

oh for real? sorry :(

OpenStudy (anonymous):

so then it should be 48/31

OpenStudy (anonymous):

it's ok no worries

OpenStudy (anonymous):

yes that's write

OpenStudy (anonymous):

thanks guys !

OpenStudy (anonymous):

how about this 1

OpenStudy (anonymous):

welcome! :)

OpenStudy (anonymous):

5 minutes to 5 hours

OpenStudy (anonymous):

5 to 15*

OpenStudy (anonymous):

5*60=300

OpenStudy (anonymous):

oops i ment 15 to 5*

OpenStudy (anonymous):

15/300

OpenStudy (anonymous):

can be reduced to 1/20

OpenStudy (anonymous):

do you have a mac?

OpenStudy (anonymous):

because if you go to the dashboard there is a whole converter thing! it is so nice! Highly recommend it